MA113 Information for instructors Spring 2002 Course coordinator: Russell Brown Textbook: The textbook for this course will be Calculus, 3rd edition, by James Stewart, ISBN 0-534-21798-2 Syllabus: A course calendar with suggested problems is available. You may duplicate the version I have provided to Elizabeth. Each instructor should also provide an individual syllabus giving name, contact information and office hours and additional information about grading policy. The course calendar and a sample syllabus are available at http://www.ms.uky.edu/∼rbrown/courses/ma113cc.s.02 as tex documents. I expect that each instructor will prepare an additional page that details grading policies and gives their name, office hours etc. The University Ombudperson has made several requests for syllabi in all courses. 1. Syllabus should describe curving policy. This is apparently in response to instructors who used “negative curves”. 2. Describe carefully what is meant by an excused absence, if attendance will be used in grading. Please see the ombud’s memo for the full story. Math resource center: The mathematics resource center or Mathskeller opened last December in room 65 in the basement of the Classroom Building. All teaching assistants for MA113 are asked to schedule at least one of their office hours in this facility (and a total of three office hours). When the schedule is fixed, it will be circulated to all students in MA113 and 114. It is hoped that the extra work involved in holding office hours outside of the department will be balanced by having a quieter offices to work in the rest of the day. Pretest: Some (but not all) of our problems in calculus are due to the poor preparation of our students. I would like to try to administer KEMTP to all students in their first recitation. More details will follow. As an alternative, you may ask that students take this test online and there is mechanism for you to collect their score. Exams: There will be three exams and a final. These exams are scheduled in the evening at the times indicated in the common exam schedule. Rooms for exams will be assigned after classes begin. I would like to suggest that the instructors for MA113 prepare common exams for the day sections. This will probably work best if common homework assignments are made in all classes so as to minimize surprises at exam time. Please emphasize to your students that they should try to clear their schedule for these exams. Students are allowed to register for a course that meets at the same time as a common exam. If an exam and course conflict, then the instructor is required to give an alternate exam. See p. 19 of the printed schedule book for details. Instructors in evening classes generally give their exams during a regularly scheduled class meeting.

Homework: I expect that we will have undergraduate paper graders for MA113. The problems in the syllabus are a guide to the topics to be covered. Assignments to be graded should be taken from nearby even numbered problems. The quality of graders is uneven. You should expect to spend some time providing guidance to your grader. MA193: In addition, to the 4 hours of credit for MA113, the department offers one additional hour of credit as MA193 on a pass/fail basis. Instructors may set their own grading policy for MA193. I suggest that students be passed in MA193 if they pass calculus and if they have no more than two unexcused absences. Thus, a student with three or more unexcused absences will fail. Below are a few commons questions about MA193. Must a student take MA193? No, unless the student is in MathExcel. MathExcel students must take MA193. What section of MA193 should a student register in? Students should register in same section number for both MA193 and MA113. If a student drops or changes sections of MA113, they should also drop or change sections of MA193. Instructors and teaching assistants should check their MA193 rolls near the end of the semester. If there is a name that is unfamiliar, please try to determine if the student is registered in another section of MA113. Can MA193 be repeated? Yes, though there is little benefit to this. Can MA193 be taken without MA113? Yes, though it is not recommended. If the suggested grading is used, such students will fail MA193. Suggested grading for MA113: Students need 90% for an A, 80% for a B, 70% for a C and 60% for a D. Grades should be based on three hour exams, quizzes, homework, and longer assignments.

3 midterm exams  300
Final exam  100
Homework, quizzes, etc. 150
TOTAL  550

Problems: Students should try to resolve problems with their instructor and/or teaching assistant. If this is not unsuccessful, the departmental ombudperson, Ted Suffridge, or the university ombudperson. If you have complaints about the syllabus, please let me know. We will need a new course coordinator next fall. Calculators: Most students are familiar with graphing calculators such as the TI-82. These calculators allow students to graph functions, solve equations, evaluate derivatives and definite integrals numerically. Elizabeth has TI-82 calculators that each instructor may check out. I suggest that students be allowed to use such calculators on exams. Test questions should be written so that it is clear whether a numerical answer from the calculator is acceptable or if students must carry out the computation by hand. In addition, some students will have machines that can carry out symbolic computations. I suggest that students not be allowed to use such machines on exams. Examples of such machines include the TI-89, TI-92, the HP48 and, of course, laptop computers. Computer labs: The math department has a computer lab, Inslab, that instructors may use. See http://www.ms.uky.edu/∼inslab for more information. This lab have Maple and Matlab available. The university also maintains numerous computer labs where students may use Maple and other mathematical software. These labs contain classrooms that may be reserved.

Russell Brown January 2, 2002